# User Location Disclosure Fails to Deter Overseas Criticism but Amplifies Regional Divisions on Chinese Social Media

### v1

@@ 4.3 Location disclosure deepened regional divisions and reduced participation on local issues
Why did domestic users, particularly those commenting on regions other than their own, become less engaged in discussions of local issues after the introduction of IP location disclosure? Our analysis suggests that the policy reshaped both the interpretation and perceived risks of online expression. By making users’ provincial origins visible, it increased the salience of regional identity and transformed comment sections into sites of heightened intergroup sensitivity. This change not only altered the tone of discourse but also deterred participation—especially from out-of-province users.
@@
Previously, critical comments on local issues could appear anonymous or broadly directed. After location disclosure, they became traceable to specific locations. As a result, even generic remarks were easily interpreted as attacks from outsiders. Figure 29(b) in the SM provides an example. A Fujian-based user, commenting on population decline in the northeast, invokes the stereotype that “mafia is everywhere” in Heilongjiang, Liaoning, and Jilin. The visible location tag invites retaliation: users target Fujian in return, citing regional tropes about fraud and scams. What might have been a passing remark now escalates into interprovincial antagonism.
@@
To assess whether this dynamic was isolated or systemic, we examine second-tier replies across posts using an LLM to detect regionally discriminatory language. Figure 3(a) shows a sharp rise in such language after location disclosure in the replies to front-page comments. Moreover, as shown in Figure S26 in the SM, this increase is driven almost entirely by cross-province exchanges; within-province interactions remain stable. These patterns indicate that location disclosure deepened existing regional divisions, raising the social and reputational cost of participating in public discussions about other regions.
This dynamic helps explain the sharp decline in comment activity on posts about local issues discussed earlier. As users became more aware of each other’s regional identity, the risk of being targeted with regionally discriminatory language increased—not only for those posting comments but also for those considering whether to reply. When even neutral or critical remarks could trigger interprovincial backlash, users likely became more hesitant to participate, especially in discussions involving other provinces. These conflicts reduced the likelihood that comments would attract replies, even when they remained visible. Figure 3(b) reflects this shift: after the introduction of location disclosure, a smaller share of comments received responses, indicating thinner and more cautious engagement on the social media platform.
@@
Moreover, we quantify the decline in user interaction by measuring the rotation among first-page comments using the CoV of comment floor numbers—a proxy for the dynamism of the comment section. Figure 3(c) shows a clear drop in CoV following the implementation of the location disclosure policy. As location tags heightened the risk of regional backlash, users became more reluctant to engage, especially on posts involving other provinces. This led to slower turnover among top comments and a more static comment hierarchy, consistent with a broader chilling effect on interaction.
@@
In sum, user location disclosure reconfigured both the meaning and the risks of online discourse. By making regional identity salient, it raised the cost of participation—especially across provincial boundaries—leading to reduced comment volume, lower dissent, and thinner engagement.
@@ 4.4 Robustness checks
We conduct a series of robustness checks, detailed in the SM, to verify the stability of our main findings on the decline in comment volume due to the implementation of the user location disclosure policy. First, we vary the bandwidths used in the local smoothing, including those selected by the rdrobust package, bandwidths chosen via cross-validation, and a range of fixed windows (8, 24, 72, 96, and 120 hours). In all cases, the observed decline in comment volume remains consistent. Second, we re-estimate the models using a log-transformed outcome, $\log(\#comments+1)$, to assess sensitivity to outliers. While the main analyses use the number of unique comments to accommodate posts with zero comments, the log transformation yields qualitatively similar results. Finally, we perform placebo tests by shifting the timing of the policy intervention to alternative dates. These placebo specifications show no discontinuities, reinforcing that the observed effects are indeed tied to the actual timing of the location disclosure policy.
@@ 5 Conclusion
We investigate the impact of a user location disclosure policy on Sina Weibo, a prominent social media platform in China. Although the policy seemed aimed at deterring engagement by overseas users, it had little impact on their actual participation. Instead, it suppressed domestic users’ engagement, particularly in commenting on out-of-province issues. The decline appears to be driven by heightened exposure to regional discrimination, as location tags made users more vulnerable to backlash when commenting on issues beyond their home provinces.
Our analyses show that authoritarian regimes can suppress criticism and shape online discourse by leveraging existing social divisions. Following the location disclosure mandate on Weibo, regional tensions among users intensified the constraints on expression. As conversations about local affairs grow more insular and uniform, the regime may become less able to identify emerging discontent—undermining its responsiveness and the long-term quality of governance (Chen and Xu, 2017b).
Supplementary Material
@@
Figure 2(a) Topic: International #Comments $32$ 0.369 107 0.342 0.545 0.407 0.396 0.435 0.454
@@
Figure 2(a) Topic: Non-International #Comments $32$0.011 93 -0.010 0.214 0.005 0.004 0.019 0.027
@@
Figure 2(b) Topic: International #Comments $96$ 0.407 122 -2.078 0.378 -1.783 -0.191 0.407 0.661
@@
Figure 2(b) Topic: Non-International #Comments $16$4.110 95 -4.603 -0.019 -4.607 -3.711 -3.353 -3.178
@@
Figure 2(c) Topic: Non-Local #Comments $40$1.399 89 -1.473 1.201 -1.511 -0.631 -0.194 0.030
@@
Figure 2(c) Topic: Local #Comments $32$7.794 99 -7.793 -0.887 -7.599 -6.748 -6.392 -6.224
@@
Figure 2(e) User: In-Province #Comments $48$1.710 88 -1.896 -0.121 -1.869 -1.500 -1.426 -1.399
@@
Figure 2(e) User: Out-of-Province #Comments $40$5.665 103 -5.702 -0.801 -5.559 -5.008 -4.735 -4.600
@@
Figure 2(e) User: Oversea #Comments $96$0.073 96 -0.108 0.060 -0.100 -0.086 -0.073 -0.066
@@
Figure 2(f) User: In-Province #Critical Comments $40$0.572 88 -0.615 -0.436 -0.596 -0.410 -0.370 -0.355
@@
Figure 2(f) User: Out-of-Province #Critical Comments $88$3.591 38 -8.153 -7.879 -6.730 -4.181 -3.430 -3.195
@@
Figure 2(f) User: Oversea #Critical Comments $96$0.030 81 0.004 0.026 -0.001 -0.036 -0.030 -0.025

### v2

@@ 4.3 Location disclosure deepened regional divisions and reduced participation on local issues
To explain why domestic users—particularly those commenting on regions other than their own—became less engaged in local discussions following the introduction of IP location disclosure, we consider three possible mechanisms.
The first, and perhaps most intuitive, explanation is a direct chilling effect. The presence of a location tag may have reminded users that their activity was being monitored by the state, thereby discouraging critical commentary. However, this explanation is unlikely to be the primary driver. As shown in Figures 2(b) (blue line) and 2(c) (red line), domestic users did not reduce their engagement on international and national (non-local) issues, including posts that attracted criticism.
A second possibility is that location tags undermined the perceived credibility of out-of-province commenters by signaling their lack of local knowledge. This interpretation aligns with the policy’s stated goal of curbing misinformation—perhaps charitably understood as discouraging speech from those unfamiliar with local context. If this mechanism were driving the observed decline, however, we would expect similar reductions in two cases: (a) overseas users commenting on domestic topics, and (b) domestic users commenting on international topics, as both groups are equally or more likely to lack firsthand knowledge. Yet we observe no such decline (Figure 2(a), red line; Figure 2(b), blue line), suggesting that concerns about credibility alone cannot explain the pattern.
A third possibility, supported by widespread anecdotal evidence, is that domestic users became more hesitant to comment on out-of-province issues due to heightened risk of backlash from local users. By making users’ provincial origins visible, the policy increased the salience of regional identity and turned comment sections into sites of intergroup tension. This mechanism should not affect comments on posts about international or national (nonlocal) issues, where location tags are less likely to provoke region-specific pushback.
@@
Before location disclosure, critical remarks on local issues often appeared anonymous or broadly directed. Afterward, they became traceable to specific locations. Even generic observations could now be read as targeted attacks from outsiders. Figure 29(b) in the SI illustrates this dynamic: a user from Fujian, commenting on population decline in the northeast, draws on the stereotype that “mafia is everywhere” in Heilongjiang, Liaoning, and Jilin. The visible location tag triggers retaliatory responses that invoke tropes about fraud and scams in Fujian. What might have been a passing comment escalates into an interprovincial confrontation.
@@
To assess whether this pattern was isolated or systemic, we use an LLM to detect regionally discriminatory language in second-tier replies. Figure 3(a) shows a sharp rise in such language following the introduction of location tags. As shown in Figure S26 in the SI, this increase is almost entirely driven by cross-provincial interactions; within-province exchanges remain stable. These findings suggest that the policy deepened existing regional divisions and raised the social and reputational costs of engaging in discussions about other regions.
This dynamic helps explain the observed decline in comment activity on local posts. As regional identity became more visible, users faced greater risks of backlash—not just for commenting critically, but for engaging at all. Even neutral remarks could provoke confrontation, discouraging participation and reducing the likelihood that comments would receive replies. Figure 3(b) captures this shift: after location disclosure, a smaller share of comments attracted responses, indicating thinner, more cautious engagement.
@@
We further quantify this decline by measuring the rotation among first-page comments using the coefficient of variation (CoV) of comment floor numbers—a proxy for dynamism in the comment section. Figure 3(c) shows a clear drop in CoV following the policy rollout. As users became more hesitant to comment, especially on posts involving other provinces, top comments were less frequently displaced, leading to a more static hierarchy of engagement.
@@
In sum, user location disclosure reconfigured both the meaning and the risks of online discourse. By making regional identity visible, it raised the cost of participation—especially across provincial boundaries—resulting in lower comment volume, diminished dissent, and a narrower, more fragmented public conversation.
@@ 4.4 Robustness checks
We conduct a series of robustness checks, detailed in the SI, to verify the stability of our main findings on the decline in comment volume due to the implementation of the user location disclosure policy. First, we vary the bandwidths used in the local smoothing, including those selected by the rdrobust package, bandwidths chosen via cross-validation, and a range of fixed windows (8, 24, 72, 96, and 120 hours). In all cases, the observed decline in comment volume remains consistent. Second, we re-estimate the models using a log-transformed outcome, $\log(\#comments+1)$, to assess sensitivity to outliers. While the main analyses use the number of unique comments to accommodate posts with zero comments, the log transformation yields qualitatively similar results. Finally, we perform placebo tests by shifting the timing of the policy intervention to alternative dates. These placebo specifications show no discontinuities, reinforcing that the observed effects are indeed tied to the actual timing of the location disclosure policy.
@@ 5 Conclusion
We study the impact of a user location disclosure policy on Sina Weibo, focusing on how the introduction of IP-based location tags affected online engagement. Although framed as a way to curb misinformation and foreign influence, the policy had little effect on overseas participation. Instead, it sharply reduced domestic users’ critical engagement with out-of-province issues. It appears that this decline was driven not by fear of state repression but by increased exposure to regionally discriminatory backlash.
These findings show that authoritarian regimes can shape discourse not only through direct censorship but also by enabling peer-based sanctions via platform design. Making identity visible can unintentionally activate social cleavages that suppress participation and narrow the scope of expression. While this may help stabilize discourse in the short run, it risks muting early signals of discontent, with potential costs for the quality of governance (Chen and Xu, 2017b).
More broadly, our study contributes to a growing literature on how platform architecture shapes user behavior. Research on algorithmic curation, identity labeling, and credibility signals shows that seemingly minor design choices can affect exposure, trust, and sharing (Guess et al., 2023; Liang et al., 2022; Nassetta and Gross, 2020; Papakyriakopoulos et al., 2022; Bradshaw et al., 2024). We extend this work by demonstrating that identity cues alone can alter participation even without content removal or fact correction. This suggests that platform design decisions—whether in democracies or authoritarian states—can shape the boundaries of public discourse in subtle, unintended, but consequential ways.
Supplementary Information
@@
Figure 2(a) Topic: International #Comments 323232 0.369 107 0.342 0.545 0.407 0.396 0.435 0.454
@@
Figure 2(a) Topic: Non-International #Comments 323232 -0.011 93 -0.010 0.214 0.005 0.004 0.019 0.027
@@
Figure 2(b) Topic: International #Comments 969696 0.407 122 -2.078 0.378 -1.783 -0.191 0.407 0.661
@@
Figure 2(b) Topic: Non-International #Comments 161616 -4.110 95 -4.603 -0.019 -4.607 -3.711 -3.353 -3.178
@@
Figure 2(c) Topic: Non-Local #Comments 404040 -1.399 89 -1.473 1.201 -1.511 -0.631 -0.194 0.030
@@
Figure 2(c) Topic: Local #Comments 323232 -7.794 99 -7.793 -0.887 -7.599 -6.748 -6.392 -6.224
@@
Figure 2(e) User: In-Province #Comments 484848 -1.710 88 -1.896 -0.121 -1.869 -1.500 -1.426 -1.399
@@
Figure 2(e) User: Out-of-Province #Comments 404040 -5.665 103 -5.702 -0.801 -5.559 -5.008 -4.735 -4.600
@@
Figure 2(e) User: Oversea #Comments 969696 -0.073 96 -0.108 0.060 -0.100 -0.086 -0.073 -0.066
@@
Figure 2(f) User: In-Province #Critical Comments 404040 -0.572 88 -0.615 -0.436 -0.596 -0.410 -0.370 -0.355
@@
Figure 2(f) User: Out-of-Province #Critical Comments $8$3.591 38 -8.153 -7.879 -6.730 -4.181 -3.430 -3.195
@@
Figure 2(f) User: Oversea #Critical Comments 969696 -0.030 81 0.004 0.026 -0.001 -0.036 -0.030 -0.025